# More than Memes: A Multimodal Topic Modeling Approach to Conspiracy Theories on Telegram

### v2

@@ Abstract
To address the increasing prevalence of (audio-)visual data on social media, and to capture the evolving and dynamic nature of this communication, researchers have begun to explore the potential of unsupervised approaches for analyzing multimodal online content. However, existing research often neglects visual content beyond memes, and in addition lacks methods to compare topic models across modalities. Our study addresses these gaps by applying multimodal topic modeling for analyzing conspiracy theories in German-language Telegram channels. We use BERTopic with CLIP for the analysis of textual and visual data in a corpus of ${\sim}40,000$ Telegram messages posted in October 2023 in 571 German-language Telegram channels known for disseminating conspiracy theories. Through this dataset, we provide insights into unimodal and multimodal topic models by analyzing symmetry and intersections of topics across modalities. We demonstrate the variety of textual and visual content shared in the channels discovered through the topic modeling, and propose a conceptual framework for the analysis of textual and visual discursive strategies in the communication of conspiracy theories. We apply the framework in a case study of the topic group Israel Gaza.
Disclaimer This paper shows examples of conspiracist and antisemitic content that may disturb some readers. They are included to deepen understanding of the issue and raise awareness.
@@ Introduction
The rise of social media has enabled the dissemination of harmful online content, including conspiracy theories (CTs), at an unprecedented speed and scale, and fostered the formation of communities around them (Bessi et al. 2016; Cinelli et al. 2021). Despite the increasingly multimodal nature of social media, research on CTs has predominantly focused on textual data. According to Mahl, Schäfer, and Zeng (2022), 78.5% of studies in this area analyze text, while less than 10% examine (audio-)visual data. Unsupervised methods offer an opportunity to develop flexible approaches to analyze evolving online content. They can be leveraged to explore large amounts of unstructured data, without the need of human-labeled training corpora. A small number of recent studies explored the potential of such approaches for clustering and analyzing multimodal online communication (Zannettou et al. 2018, 2019; Qu et al. 2023; Theisen and Scheirer 2024). However, existing research often neglects visual content beyond memes, and in addition lacks methods to compare topic models across modalities. Therefore, many questions remain: What information does each modality convey? How do different modalities interact? What are the specific contributions of each modality? We contribute to this field of research by exploring multimodal topic modeling of text-based, image-based, and text-image-based communication in German-language conspiracist Telegram channels. [^1] We use the term ‘multimodal’ to refer to text-image-data, although it can encompass a variety of modalities. We apply BERTopic (Grootendorst 2022) with the vision language model CLIP (Radford et al. 2021) to text, image, and text-image data. We analyze data from October 2023, a period marked by the Hamas’ attack on Israel and the Israel-Gaza war, which represent crisis-like events that often spur the proliferation of CTs (van Prooijen and Douglas 2017). Using this dataset, we offer insights into both unimodal and multimodal topic models, highlighting the diverse textual and visual content shared in the channels as revealed through the topic modeling process. Further, we aim to understand whether there is correspondence or overlap of similar topics across modalities.
@@
Visuals play an important role for the communication of CTs: images appear to provide immediate ‘knowledge’, and are powerful means in presenting complex narratives (Caumanns and Önnerfors 2020). A small number of studies examine the relevance of visual content in CTs: Tuters and Willaert (2022) analyze over 470,000 CT-related Instagram posts, guided by the concept of narrative convergence, showing that different conspiracy narratives (e.g., anti-vax, QAnon, and anti-5G), converge over time, with figures like Donald Trump and Bill Gates playing central roles in connecting them. Holzer (2021) finds that visual communication triggers more activity than text, and is strategically used to circumvent detection mechanisms, or to tactically disguise hateful content as humour using emojis or memes (Al-Rawi 2024). Qualitative research of CT-related memes has shown that they serve to build ingroup community and criticize outgroups like governments and elites (Gür-Şeker, Boonen, and Wentker 2022). However, other research on far-right visual communication on social media reveals that memes are less prevalent compared to other image types such as photographs and promotional materials (McSwiney et al. 2021), or screenshots of news and social media posts (Sosa and Sharoff 2022). This indicates an important gap in existing research with its focus on memes. Clustering and topic modeling (and combinations thereof) are widely used approaches for partitioning data into meaningful groups. Depending on the underlying mathematical approach, they not only assign documents to one or more groups but might also allow to discover latent themes within large corpora, each representing a “human-interpretable semantic concept” (Abdelrazek et al. 2023). Respective methods involve technical steps such as feature extraction, typically via pre-trained models, and the application of clustering algorithms, but also human-guided steps such as validation and interpretation of results (Peng, Lock, and Ali Salah 2023). While these approaches allow for the discovery of textual, visual or multimodal ‘topics’, their application to social and political phenomena comes with the challenge of connecting automatically generated clusters with theoretically meaningful concepts post-hoc, a process which requires background knowledge and cultural understanding (Peng, Lock, and Ali Salah 2023). The choice of modality is relevant to the technical steps (i.e. how the data is processed, what features are extracted), but is also likely to influence the human interpretation of the topics.
@@
Conspiracy ideology 12,177 (30.4%) 186
QAnon 11,894 (29.7%) 73
Esotericism 5,623 (14.0%) 89
Querdenken 4,508 (11.2%) 125
COVID-19 Disinformation 3,538 (8.8%)) 72
Reichsbürger 2,370 (5.9%) 26
@@
Table 1: Channel category distribution in the base corpus.
@@
The validation and interpretation of topic modeling results can be challenging, and available quantitative measures do not necessarily correspond with human judgment. Manual inspection of topics and the assignment of names (topic assignment) is a common validation approach (Zhang and Peng 2022; Ying, Montgomery, and Stewart 2022). In the context of our study, topic annotations were conducted by an expert annotator with field knowledge and practical experience in social media data annotation. In addition to the assignment of a topic name, the topic type and image type of each topic were annotated. The corresponding label sets were developed inductively after an initial exploration of the results.
@@
For validation, the same annotator re-annotated the topic type and the image type assignments on a random sample of 50 topics per modality after three days. The intra-annotator agreement, measured as Cohen’s $\kappa$ (see Table 3), was moderate to high, except for the topic type in the image-only setting. Error analysis revealed that inconsistencies in annotating image topics stemmed from relying on stylistic similarities rather than semantic coherence (Zhang and Peng 2022, cf.). We resolved label conflicts to obtain the final topic and image type labels.
@@
To improve efficiency, we evaluated the use of a generative Large Language Model (LLM), gemini-1.5-flash, for annotating and naming topics. The model was prompted to generate concise topic names and assign topic and image types (using a simplified label set for image types) and short definitions for each label. [^8] Prompt details can be found in the accompanying code repository. The model’s performance was evaluated against human annotations as the gold standard, with macro average F1 scores showing acceptable to good agreement (see Table 3). For topic name evaluation, we computed cosine similarity between manual and LLM-generated names using the paraphrase-multilingual-MiniLM-L12-v2 model. The median similarity scores indicate moderate to high similarity for all modalities (text: 0.69, image: 0.62, multimodal: 0.65). The annotation process was reduced from about 4 hours (manual) to 15 minutes per modality, while topic grouping (around 30 minutes in total), remained manual to preserve human assessment.
@@ Results
Examining and comparing topic types and images types across modalities allows us to answer RQ 1. As Table 4 shows, the topic type content is predominant across modality settings, with the highest proportion found in the image-based topic model. Interpreting image-only topics appears to be easier than text-based or multimodal topics. The higher proportion of advertising topics in November compared to October, along with the lower proportion of caption topics, may result from overlap in the content or characteristics of these two types. The higher proportion of content unclear in the multimodal setting could be due to the difficulty for human annotators to interpret dual-modality topics, which requires to consider both textual and visual aspects of a topic. The predominant image type category in October 2023 is photo, followed by screenshots, which in total account for around 20% of the topics. The prominence of screenshots increases in November 2024, reaching around 50%, while the photo proportion decreases but remains relevant. Memes range between 1% and 5% across modalities and time periods. Maps drop from 6-7% in October 2023 to (less than) 1% in November 2024, likely due to reduced relevance of geopolitical topics like Israel Gaza. The low proportions of infographics and maps also reflect differences in category granularity, with broader categories like photo and screenshots dominating. However, despite also being a broad category, memes remain minor, clearly indicating they are not key genres.
@@
The manually assigned topic groups offer a broad view of thematic differences and similarities across modalities. Table 5 shows the five largest topic groups per modality (excluding the Misc group due to its lack of thematic expressiveness) in October 2023. Across all modalities, Israel Gaza dominates, reflecting the dataset’s focus on October 2023, including the Hamas attack on Israel (Oct 7), and the ensuing war. Its prevalence varies by modality, with 36.2% in text, 12.9% in images, and 16.7% in the multimodal data. In the text-based setting, it forms a single large topic, while in the image and multimodal setting it is divided into 18 and 11 smaller topics, respectively. Text captures more messages with a clear main focus but lacks the finer granularity provided by image and multimodal settings, which offer more detailed perspectives at the cost of comprehensiveness. The image and multimodal setting share the topic group Protests, while Ukraine War appears in text and image, though with differences in rank and size. The topic group Self-Love, present in text and multimodal settings, likely originates from channels in the Esotericism channel category. Among these top-ranked topic groups, we will focus on Israel Gaza, Protests, and Ukraine war in the following, as they exhibit a greater potential for CTs compared to self-love topics.
@@
In November 2024, we find three topic groups ranking in the top five across all modalities: German Politics, Ukraine War, and US Politics. COVID-19 appears in text and multimodal, while Finance, Military & War, and Conspiracy (specific) each rank in the top five of one modality. This marks a clear thematic shift, with Israel Gaza losing its dominant role.
@@
Figure 1 shows the message frequency over time for the dominant topic groups. Israel Gaza shows a sharp increase on October 7 (Hamas’ attack on Israel) and a slight decline over time. Text dominates with about 50% of messages, while image and multimodal make up roughly 25% each. We found consistent modality-based proportions over time for the other (smaller) topic groups as well, with differing modality distributions. E.g., Ukraine war is mostly represented through multimodal topics (two-thirds) and text (one-third), with marginal relevance of image topics. In contrast, Protests is dominated by image (two-thirds), and multimodal topics (one-third), with minimal text presence.
@@
The topic intersections across models differ notably between the dominant topic groups in October 2023 (see Figure 2). Israel Gaza has an overlap of all three settings, but also distinct documents unique to each. Around 60% of documents in the text group are distinct, while image and multimodal show larger overlap. In contrast, the Protests topic group has a very small core, but a large overlap of text and multimodal. The Ukraine War topic group completely misses a shared core, and generally shows little overlap. Almost 90% of the documents from the multimodal setting are not identified by the text nor the image model. While Israel Gaza is predominantly text-based, Protests is dominated by the image setting. For Ukraine War, multimodal messages prevail, and the image setting is marginal. We found similar variations of overlap in the November 2024 data. These insights show that topic representation varies not only by input modality but also by topic. This requires further investigation to understand what drives modality predominance per topic, or how core documents differ from those found in only one modality setting of a topic. After this quantitative investigating of topics, we now turn to qualitative insights into explore CT-related discursive strategies in greater depth.
@@
Previous studies provide important insights into textual CT-related strategies. The Rumor Interaction Analysis System (RIAS), originally developed by Bordia and Difonzo (2004), for analyzing rumor discussions in newsgroups, categorizes strategies like authenticating claims (e.g., citing media or personal expertise), providing information, directives, and using rhetorical or sarcastic statements. Their study shows how people attempt to persuade each other, often citing personal experiences or media sources to confirm or contradict rumors. The five most common types of statements are interrogatory statements, sense-making (attempting to falsify or verify a rumor), providing information, disbelief, and digressive (unrelated) statements. The authors conceptualize rumors as efforts to reduce uncertainty in crisis-like, often ambiguous situations, typically involving events with significant impact – characteristics shared with CTs. These commonalities motivated Wood (2018) to adapt the RIAS schema for analyzing CTs related to the Zika virus on Twitter in 2015-2016, focusing on six categories: reference, belief, disbelief, authenticating, directives, and rhetorical questions. They find that CT-supporting tweets more often cite authorities and use rhetorical questions compared to CT-opposing posts. We incorporate authenticating, directives, and rhetorical questions into our framework, as the other categories do not represent strategies, but refer to stance, which is outside the scope of our analysis. While the work of Wood (2018) offers some suitable categories, we find the range of provided strategies too limited for analyzing our data. To extend the set of strategies, the work of Kou et al. (2017) offers a valuable source. Their work draws on a different strand of rumor research (a study on AIDS denialism byMeylakhs et al. (2014)), which they adapt to analyze CT-related posts about the Zika virus on Reddit. They identify a number of strategies in CT-supporting posts including selective citation of authoritative sources, cultural references (e.g., movies), narrative elaboration (adding unverifiable details), ‘connecting the dots’ (linking unrelated information), ideological arguments (e.g., system critique), casting doubt, and promoting skepticism. While selective citation and connecting the dots align with authentication in Wood (2018), other strategies such as system critique, casting doubt, and promoting skepticism demonstrate how conspiracy theorists undermine official narratives while constructing alternative explanations. We consider this strategy of discrediting mainstream accounts as important aspect of CTs, and therefore include it in our framework.
@@
We used the proposed framework to analyze the predominant topic group in the October 2023 dataset, Israel Gaza. For each modality, we explored the thematic aspects of the topics and analyzed the textual and visual discursive strategies used in CT communication. [^9] Note that all text examples are presented as English translations of the posts.
@@
The identified topics mention political and military actors (individuals and groups) in the region, and cover different aspects of military actions in Gaza, with the attack on a hospital in Gaza standing out as a distinct event. They also refer to protests against Israel’s actions, and discuss the role of the US administration and the German government in the war. Other topics address the history of the Israeli-Palestinian conflict, or claim to reveal the reputed ‘truth’ about the Hamas attack on Israel. More general topics include Zionism, media propaganda, and the purported relation between Islam, immigration, and antisemitism.
Applying the conceptual framework revealed a number of discursive strategies: Topics discussing the history of the Israeli-Palestinian conflict as well as claims to ‘reveal the truth’ about the Hamas attack can be interpreted as providing evidence. A closely related strategy are acts of disclosure and revealing connections: For instance, the topic Truth Hamas attack suggests that the attack was actually orchestrated by Israel, the USA, ‘Zionists’, and Prime Minister Netanyahu in collaboration with a ‘globalist elite’. This claim explicitly connects multiple actors and aligns them with a shared, malicious goal, reinforcing a conspiratorial worldview. One particularly explicit post claims that the incident was “an ‘inside job’ by the globalist elite” in cooperation with the USA and Israel, “as part of the grand master plan for World War III.” The strategy of authentication occurs in posts which claim to refer to insider knowledge (without naming them explicitly). For example, a user refers to unnamed “insiders” and “officials”, which creates an impression of credibility. Irony and sarcasm are employed to discredit mainstream accounts: The following message sarcastically criticizes the framing of oppositional views in mainstream discourse: “Corona deniers, anti-vaccinationists and Putin haters are out! Now there are ISRAEL HATERS.” The statement also reflects the ingroup vs. outgroup strategy, portraying the ingroup as marginalized rebels unjustly dismissed by the mainstream. Another post adapts Reichsbürger rhetoric to a new context by claiming that Israel is not a nation-state but a corporate entity. This aligns with the strategy of constructing enemy images by using delegitimizing language. An example of emotionalization is this post, which uses capslock and exaggerations while also serving as an act of disclosure, claiming the true goal behind the Israel-Gaza war is:
@@
The image-based topics feature diverse visuals, including photos of explosions and debris, portraits of key figures (e.g., Benjamin Netanyahu), screenshots, memes, graphic representations like maps, cartoons, and symbolic images such as flags. Like in the text setting, topics cover political and military actors, aspects of the Israel Gaza war, and the history of the conflict. Photos of explosions and debris (Figure 3, row 1), highlighting the war’s devastation and evoking fear, exemplify emotionalization. Symbols like flags (e.g., of Ukraine and Israel, Figure 3, row 2), visually reveal connections, linking nations or ideologies to construct associations. Memes and cartoons (Figure 3, row 2) support the strategy of opposing or discrediting mainstream accounts. These visuals often employ irony or sarcasm and create links between unrelated events such as the Israel-Gaza war, the Ukraine war, and COVID-19. In doing so, they reinforce oppositional narratives through humor and simplification. Textual elements play a central role in many images, with screenshots being the most text-heavy examples. The technique of ‘quoting’ via screenshots, where text from other sources is directly captured and presented, aligns with strategies of providing evidence and authenticating. Highlighting details in screenshots serves to emphasize key information, expose inconsistencies, or delegitimize specific actors (Figure 3, row 3). Maps and infographics often serve the purpose of providing evidence but can also reflect the ideological framing. For example, some maps label the Hamas attack as ‘Palestinian uprising’, implicitly framing Israel as a suppressor and supporting the construction of enemy images (Figure 3, row 4).

### v3

@@ Abstract
[…]
@@ Introduction
The rise of social media has enabled the dissemination of harmful online content, including conspiracy theories (CTs), at an unprecedented speed and scale, and has promoted the formation of communities around them (Bessi et al. 2016; Cinelli et al. 2021). Despite the increasingly multimodal nature of social media, research on CTs has predominantly focused on textual data. According to Mahl, Schäfer, and Zeng (2022), 78.5% of the studies in this area analyze text, while less than 10% examine (audio)visual data. Unsupervised methods offer an opportunity to develop flexible approaches to analyze evolving online content. They can be leveraged to explore large amounts of unstructured data, without the need for human-labeled training corpora. A small number of recent studies explores the potential of such approaches for clustering and analyzing multimodal online communication (Zannettou et al. 2018, 2020; Qu et al. 2023; Theisen and Scheirer 2024b). However, existing research often neglects visual content beyond memes and, in addition, lacks methods to compare topic models across modalities. Therefore, many questions remain: What information does each modality convey? How do different modalities interact? What are the specific contributions of each modality? We contribute to this field of research by exploring multimodal topic modeling of text-based, image-based, and text-image-based communication on German-language conspiracist Telegram channels. [^1] We use the term ‘multimodal’ to refer to text-image-data, although it can encompass a variety of modalities. We apply BERTopic (Grootendorst 2022) with the vision language model CLIP (Radford et al. 2021) to text, image, and text-image data. We analyze data from October 2023, a period marked by the Hamas attack on Israel and the Israel-Gaza war, which represent crisis-like events that often spur the proliferation of CTs (van Prooijen and Douglas 2017). Using this data set, we offer insights into both unimodal and multimodal topic models, highlighting the diverse textual and visual content shared in the channels as revealed through the topic modeling process. Further, we aim to understand whether there is correspondence or overlap of similar topics across modalities.
@@
Visuals play an important role for the communication of CTs: images appear to provide immediate ‘knowledge’, and are powerful means in presenting complex narratives (Caumanns and Önnerfors 2020). A small number of studies examine the relevance of visual content in CTs: Tuters and Willaert (2022) analyze over 470,000 CT-related Instagram posts, guided by the concept of narrative convergence, showing that different conspiracy narratives (e.g., anti-vax, QAnon, and anti-5G), converge over time, with figures like Donald Trump and Bill Gates playing central roles in connecting them. Holzer (2021) finds that visual communication triggers more activity than text, and is strategically used to circumvent detection mechanisms, or, as Al-Rawi (2024) shows, to tactically disguise hateful content as humor using emojis or memes. Qualitative research of CT-related memes has shown that they serve to build ingroup community and criticize outgroups like governments and elites (Gür-Şeker, Boonen, and Wentker 2022). However, other research on far-right visual communication on social media reveals that memes are less prevalent compared to other image types such as photographs and promotional materials (McSwiney et al. 2021), or screenshots of news and social media posts (Sosa and Sharoff 2022). This indicates an important gap in existing research with its focus on memes.
Clustering and topic modeling (and combinations thereof) are widely used approaches for partitioning data into meaningful groups. Depending on the underlying mathematical approach, they not only assign documents to one or more groups, but might also allow to discover latent themes within large corpora, each representing a “human-interpretable semantic concept” (Abdelrazek et al. 2023). Respective methods involve technical steps such as feature extraction, typically via pre-trained models, and the application of clustering algorithms, but also human-guided steps such as validation and interpretation of results (Peng, Lock, and Ali Salah 2023). While these approaches allow for the discovery of textual, visual or multimodal ‘topics’, their application to social and political phenomena comes with the challenge of connecting automatically generated clusters with theoretically meaningful concepts post-hoc, a process which requires background knowledge and cultural understanding (Peng, Lock, and Ali Salah 2023). The choice of modality is relevant to the technical steps (i.e. how the data is processed, what features are extracted), but is also likely to influence the human interpretation of the topics.
@@
Conspiracy Ideology 12,177 (30%) 186
QAnon 11,894 (30%) 73
Esotericism 5,623 (14%) 89
Querdenken 4,508 (11%) 125
COVID-19 Disinfo 3,538 (9%) 72
Reichsbürger 2,370 (6%) 26
@@
Table 1: Channel category distribution in the base corpus (percentages rounded).
@@
The validation and interpretation of topic modeling results can be challenging, and available quantitative measures do not necessarily correspond with human judgment. Manual inspection of topics and the assignment of names (topic assignment) is a common validation approach (Zhang and Peng 2022; Ying, Montgomery, and Stewart 2022). In the context of our study, topic annotations were conducted by the author, who has field knowledge and practical experience in social media data annotation. In addition to the assignment of a topic name, the topic type and image type of each topic were annotated. The corresponding label sets were developed inductively after an initial exploration of the results.
@@
For validation, the author re-annotated the topic type and the image type assignments on a random sample of 50 topics per modality after three days. The intra-annotator agreement, measured as Cohen’s $\kappa$ (see Table 3), was moderate to high, except for the topic type in the image-only setting. Error analysis revealed that inconsistencies in annotating image topics stemmed from relying on stylistic similarities rather than semantic coherence (Zhang and Peng 2022, cf.). We resolved label conflicts to obtain the final topic and image type labels.
@@
To improve efficiency, we evaluated the use of a generative Large Language Model (LLM), gemini-1.5-flash, for annotating and naming topics. The model was prompted to generate concise topic names and assign topic and image types (using a simplified label set for image types) and short definitions for each label. [^8] For prompt details please refer to the accompanying code repository. The model’s performance was evaluated against human annotations as the gold standard, with macro average F1 scores showing acceptable to good agreement (see Table 3). For topic name evaluation, we computed cosine similarity between manual and LLM-generated names using the paraphrase-multilingual-MiniLM-L12-v2 model. The median similarity scores indicate moderate to high similarity for all modalities (text: 0.69, image: 0.62, multimodal: 0.65). The annotation process was reduced from about 4 hours (manual) to 15 minutes per modality, while topic grouping (around 30 minutes in total), remained manual to preserve human assessment.
@@ Results
Examining and comparing topic types and images types across modalities allows us to answer RQ 1. As Table 4 shows, the topic type content is predominant across modality settings, with the highest proportion found in the image-based topic model. Interpreting image-only topics appears to be easier than text-based or multimodal topics. The higher proportion of advertising topics in November compared to October, along with the lower proportion of caption topics, may result from overlap in the content or characteristics of these two types. The higher proportion of content unclear in the multimodal setting could be due to the difficulty for human annotators to interpret dual-modality topics, which requires to consider both textual and visual aspects of a topic.
The predominant image type category in October 2023 is photo, followed by screenshots, which in total account for around 20% of the topics. The prominence of screenshots increases in November 2024, reaching around 50%, while the photo proportion decreases but remains relevant. Memes range between 1% and 5% across modalities and time periods. Maps drop from 6-7% in October 2023 to (less than) 1% in November 2024, likely due to reduced relevance of geopolitical topics like Israel Gaza. The low proportions of infographics and maps also reflect differences in category granularity, with broader categories like photo and screenshots dominating. However, despite also being a broad category, memes remain minor, clearly indicating they are not key genres.
@@
The manually assigned topic groups offer a broad view of thematic differences and similarities across modalities. Table 5 shows the five largest topic groups per modality (excluding the Misc group due to its lack of thematic expressiveness) in October 2023. Across all modalities, Israel Gaza dominates, reflecting the data set’s focus on October 2023, including the Hamas attack on Israel (Oct 7th), and the ensuing war. Its prevalence varies by modality, with 36.2% in text, 12.9% in images, and 16.7% in the multimodal data. In the text-based setting, it forms a single large topic, while in the image and multimodal setting, it is divided into 18 and 11 smaller topics, respectively. Text captures more messages with a clear main focus, but lacks the finer granularity provided by image and multimodal settings, which offer more detailed perspectives at the cost of comprehensiveness. The image and multimodal setting share the topic group Protests, while Ukraine War appears in text and image, though with differences in rank and size. The topic group Self-Love, present in text and multimodal settings, likely originates from channels in the Esotericism channel category. Among these top-ranked topic groups, we will focus on Israel Gaza, Protests, and Ukraine war in the following, as they exhibit a greater potential for CTs compared to self-love topics.
@@
In November 2024, we find three topic groups ranking in the top five across all modalities: German Politics, Ukraine War, and U.S. Politics. COVID-19 appears in text and multimodal, while Finance, Military & War, and Conspiracy (specific) each rank in the top five of one modality. This marks a clear thematic shift, with Israel Gaza losing its dominant role.
@@
Figure 1 shows the message frequency over time for the dominant topic groups. Israel Gaza shows a sharp increase on October 7th (Hamas attack on Israel) and a slight decline over time. Text dominates with about 50% of messages, while image and multimodal make up roughly 25% each. We found consistent modality-based proportions over time for the other (smaller) topic groups as well, with differing modality distributions. E.g., Ukraine war is mostly represented through multimodal topics (two-thirds) and text (one-third), with marginal relevance of image topics. In contrast, Protests is dominated by image (two-thirds), and multimodal topics (one-third), with minimal text presence.
@@
The topic intersections across models differ notably between the dominant topic groups in October 2023 (see Figure 2). Israel Gaza has an overlap of all three settings, but also distinct documents unique to each. Around 60% of documents in the text group are distinct, while image and multimodal show larger overlap. In contrast, the Protests topic group has a very small core, but a large overlap of text and multimodal. The Ukraine War topic group completely misses a shared core, and generally shows little overlap. Almost 90% of the documents from the multimodal setting are not identified by the text nor the image model. While Israel Gaza is predominantly text-based, Protests is dominated by the image setting. For Ukraine War, multimodal messages prevail, and the image setting is marginal. We found similar variations of overlap in the November 2024 data. These insights show that topic representation varies not only by input modality but also by topic. This requires further investigation to understand what drives modality predominance per topic, or how core documents differ from those found in only one modality setting of a topic. After this quantitative investigating of topics, we now turn to qualitative insights to explore CT-related discursive strategies in greater depth.
@@
Previous studies provide important insights into textual CT-related strategies. The Rumor Interaction Analysis System (RIAS), was originally developed y Bordia and Difonzo (2004) for analyzing rumor discussions in newsgroups, categorizing strategies like authenticating claims (e.g., citing media or personal expertise), providing information, directives, and using rhetorical or sarcastic statements. Their study shows how people attempt to persuade each other, often citing personal experiences or media sources to confirm or contradict rumors. The five most common types of statements are interrogatory statements, sense-making (attempting to falsify or verify a rumor), providing information, disbelief, and digressive (unrelated) statements. The authors conceptualize rumors as efforts to reduce uncertainty in crisis-like, often ambiguous situations, typically involving events with significant impact – characteristics shared with CTs. These commonalities motivated Wood (2018) to adapt the RIAS schema for analyzing CTs related to the Zika virus on Twitter in 2015-2016, focusing on six categories: reference, belief, disbelief, authenticating, directives, and rhetorical questions. They find that CT-supporting tweets more often cite authorities and use rhetorical questions compared to CT-opposing posts. We incorporate authenticating, directives, and rhetorical questions into our framework, as the other categories do not represent strategies, but refer to stance, which is outside the scope of our analysis. While the work of Wood (2018) offers some suitable categories, we find the range of provided strategies too limited for analyzing our data. To extend the set of strategies, the work of Kou et al. (2017) offers a valuable source. Their work draws on a different strand of rumor research (a study on AIDS denialism by Meylakhs et al. (2014)), which they adapt to analyze CT-related posts about the Zika virus on Reddit. They identify a number of strategies in CT-supporting posts including selective citation of authoritative sources, cultural references (e.g., movies), narrative elaboration (adding unverifiable details), ‘connecting the dots’ (linking unrelated information), ideological arguments (e.g., system critique), casting doubt, and promoting skepticism. While selective citation and connecting the dots align with authentication in Wood (2018), other strategies such as system critique, casting doubt, and promoting skepticism demonstrate how conspiracy theorists undermine official narratives while constructing alternative explanations. We consider this strategy of discrediting mainstream accounts as important aspect of CTs, and therefore include it in our framework.
@@
We used the proposed framework to analyze the predominant topic group in the October 2023 data set, Israel Gaza. For each modality, we explore the thematic aspects of the topics and analyze CT-related textual and visual discursive strategies. [^9] Note that all text examples are presented as English translations of the posts.
@@
The identified topics mention political and military actors (individuals and groups) in the region, and cover different aspects of military actions in Gaza, with the attack on a hospital in Gaza standing out as a distinct event. They also refer to protests against Israel’s actions, and discuss the role of the U.S. administration and the German government in the war. Other topics address the history of the Israeli-Palestinian conflict, or claim to reveal the reputed ‘truth’ about the Hamas attack on Israel. More general topics include Zionism, media propaganda, and the purported relation between Islam, immigration, and antisemitism.
Applying the conceptual framework reveals a number of discursive strategies: Topics discussing the history of the Israeli-Palestinian conflict as well as claims to ‘reveal the truth’ about the Hamas attack can be interpreted as providing evidence. A closely related strategy are acts of disclosure and revealing connections: For instance, the topic Truth Hamas attack suggests that the attack was actually orchestrated by Israel, the USA, ‘Zionists’, and Prime Minister Netanyahu in collaboration with a ‘globalist elite’. This claim explicitly connects multiple actors and aligns them with a shared, malicious goal, reinforcing a conspiratorial worldview. One particularly explicit post claims that the incident was “an ‘inside job’ by the globalist elite” in cooperation with the USA and Israel, “as part of the grand master plan for World War III.” The strategy of authentication occurs in posts which claim to refer to insider knowledge (without naming them explicitly). For example, a user refers to unnamed “insiders” and “officials”, which creates an impression of credibility. Irony and sarcasm are employed to discredit mainstream accounts: The following message sarcastically criticizes the framing of oppositional views in mainstream discourse: “Corona deniers, anti-vaccinationists and Putin haters are out! Now there are ISRAEL HATERS.” The statement also reflects the ingroup vs. outgroup strategy, portraying the ingroup as marginalized rebels unjustly dismissed by the mainstream. Another post adapts Reichsbürger rhetoric to a new context by claiming that Israel is not a nation-state but a corporate entity. This aligns with the strategy of constructing enemy images by using delegitimizing language. An example of emotionalization is this post, which uses capslock and exaggerations while also serving as an act of disclosure, claiming the true goal behind the Israel-Gaza war is:
@@
The image-based topics feature diverse visuals, including photos of explosions and debris, portraits of key figures (e.g., Benjamin Netanyahu), screenshots, memes, graphic representations like maps, cartoons, and symbolic images such as flags. Like in the text setting, topics cover political and military actors, aspects of the Israel Gaza war, and the history of the conflict. Photos of explosions and debris (Figure 3, row 1), highlighting the war’s devastation and evoking fear, exemplify emotionalization. Symbols like flags (e.g., of Ukraine and Israel, Figure 3, row 2) visually reveal connections, linking nations or ideologies to construct associations. Memes and cartoons (Figure 3, row 2) support the strategy of opposing or discrediting mainstream accounts. These visuals often employ irony or sarcasm and create links between unrelated events such as the Israel-Gaza war, the Ukraine war, and COVID-19. In doing so, they reinforce oppositional narratives through humor and simplification. Textual elements play a central role in many images, with screenshots being the most text-heavy examples. The technique of ‘quoting’ via screenshots, where text from other sources is directly captured and presented, aligns with strategies of providing evidence and authenticating. Highlighting details in screenshots serves to emphasize key information, expose inconsistencies, or delegitimize specific actors (Figure 3, row 3). Maps and infographics often serve the purpose of providing evidence but can also reflect the ideological framing. For example, some maps label the Hamas attack as “Palestinian uprising”, implicitly framing Israel as a suppressor and supporting the construction of enemy images (Figure 3, row 4).